# thanks: Corresponding author

### v1

@@ Abstract
Copyright 2025 Author(s). This article is distributed under a Creative Commons Attribution-NonCommercial-NoDerivs 4.0 International (CC BY-NC-ND) License.
preprint: xxx
@@ I INTRODUCTION
Magnetic force microscopy (MFM) is a powerful technique for resolving nanoscale magnetic textures with high field sensitivity. In MFM, a magnetically coated tip mounted on an oscillating cantilever is scanned over the sample at a measurement height $z$. The phase shift $\Delta\phi$ of the oscillating cantilever is monitored to measure the interaction strength between the tip and the sample. For many applications, it is essential to obtain not only qualitative data but also quantitative data in terms of the sample’s stray field distribution using either the magnetic field $H$ (in units of A/m) or the magnetic flux density $B=\mu_{0}H$ (in units of Tesla). Such quantitative MFM (qMFM) requires calibrating the MFM system, that means finding the functional relationship that connects the effective magnetic surface charge density $\sigma_{\text{eff}}^{\text{sample}}(x,y)$ or the magnetic field $B^{\text{sample}}(x,y)$ of a thin film sample to the measured MFM signal, thereby providing spatially resolved phase shift data $\Delta\phi(x,y)$.
@@
Initially, calibrations relied on interpreting the tip as a point-like dipole or monopole interacting with the local stray field and fitting this model to reference measurements. However, such approaches exhibit significant limitations due to their strong dependence on the sample’s feature sizes Vock et al. (2010). These limitations can be overcome by determining the spatial frequency-dependent transfer function of the MFM system, known as the instrument calibration function (ICF), as demonstrated by Hug et al. Hug et al. (1998) and Schendel et al. van Schendel et al. (2000). The ICF is a function of the cantilever’s mechanical properties and of the gradient of the force exerted on the tip. This force arises from the non-local interaction between the vertical component of the spatially extended MFM tip stray field distribution $B_{z}^{\text{tip}}$ and the magnetization of the sample under test (SUT), as illustrated in Fig. 1a. The spatially extended tip stray field distribution at the sample surface thus governs the point-spread function character of the ICF.
@@
Mathematically, the measured cantilever phase shift can be described conveniently in the partial Fourier space $(x,y,z)\to(k_{x},k_{y},z)=(\mathbf{k},z)$, where $k=\sqrt{k_{x}^{2}+k_{y}^{2}}$. In this space, the convolution between the tip stray field at the sample surface and the effective magnetic surface charge density of the sample $\sigma_{\text{eff}}^{\text{sample}}$ translates into a multiplication:
$\Delta\phi(\mathbf{k},z)=$ $\text{ICF}(\mathbf{k})\cdot\sigma_{\text{eff}}^{\text{sample}}$
@@
Here, $Q$ and $C$ represent the quality factor and the stiffness of the oscillating cantilever, respectively. The lever correction function (LCF) accounts for the finite oscillation amplitude $A$ and the canting angle $\Theta$ of the cantilever.
Consequently, for a quantitative analysis of a SUT, it is essential to know both the shape and amplitude of the tip stray field distribution, as well as the scalar geometrical and mechanical parameters of the cantilever. This information allows for the calculation of the effective magnetic charge density distribution of any sample - and thus its stray field - by deconvolving the measured phase shift signal. Therefore, calibrating MFM requires determining $B_{z}^{\text{tip}}(k)$. By exploiting Eq. I, this can be achieved by measuring the phase shift distribution $\Delta\phi^{\text{ref}}$ of a reference sample with a well-known effective magnetic surface charge density $\sigma_{\text{eff}}^{\text{ref}}$. Ideally, deconvolution to obtain $B_{z}^{\text{tip}}$ from Eq. I can be described as a division in Fourier space. However, due to measurement noise, this deconvolution is ill-posed and necessitates regularization, a method used to impose additional constraints to stabilize the inversion process and reduce noise in the reconstructed image. This is conveniently performed using an inverse Wiener filter (see Ref. Wiener, 1949) with a regularization parameter $\alpha_{1}$ (Eq. 2) as discussed in the relevant IEC standard on quantitative MFM IEC (2021). Typically, $\alpha_{1}$ is determined by an L-curve criterion Hu et al. (2020). Due to regularization, the reconstructed tip stray field distribution becomes frequency-filtered. To distinguish the reconstructed $B_{z}^{\text{tip}}$ from the actual $B_{z}^{\text{tip}}$, it is often referred to as the tip transfer function (TTF).
@@
Another factor influencing the TTF is the choice of the reference sample. To date, only maze domain patterns have been tested and validated as suitable reference samples Hug et al. (1998); Hu et al. (2020); Sakar et al. (2021a). These samples are characterized by distinct domain patterns with specific domain sizes and narrow transitions between domains. Determining the underlying magnetization pattern, which is necessary for calculating the effective surface charge density, involves discriminating the MFM image into up- and down-magnetized domains followed by applying a domain wall operator. The thereof calculated effective charge density pattern provides a good approximation of the real density pattern (Fig. 1a) and is subsequently used in the deconvolution process.
@@
An MFM image of a typical maze domain sample, specifically a CoPt multilayer stack as referenced in Ref. Hu et al., 2020, is illustrated in Fig. 1b alongside a cross-section plot and its circularly averaged Fourier spectrum. The TTF for the tip used in Ref. Sakar et al., 2021b, derived from deconvolving the measured phase shift image of this CoPt multilayer, is depicted as the black curve in Fig. 2f. Although the fundamental validity of transfer function-based calibration has been demonstrated for these specific samples, its limitations - stemming from dependencies on the reference sample characteristics and measurement parameters such as image size and pixel resolution - have not yet been thoroughly investigated.
Figure 1: (a) MFM measurements represented as the convolution of the stray field from the tip with the sample magnetization. (b) MFM phase shift data for a 5.12 µm x 5.12 µm scan of a CoPt multilayer stack, displaying a maze domain pattern. (c) Cross-section of one line within the domain pattern shown in (b). (d) Corresponding wavelength density spectrum in Fourier space.
@@
Additionally, understanding the relationship between the TTF and the actual tip stray field distribution is crucial but challenging due to difficulties in directly assessing the tip stray field distribution with sufficient resolution and sensitivity. Over the past three decades, only a few attempts have been published that characterize the magnetic stray-field distribution of MFM tips using techniques such as the Hall effect Abelmann et al. (1998); Vock et al. (2014), Lorentz tomography or holography Dai et al. (2018); Neu et al. (2018); Gao et al. (2004). Recent advancements in NV microscopy (see Ref. Casola et al., 2018) have introduced new methods for highly spatially resolved characterization of magnetic stray fields.
@@
Exploiting these techniques, a recent study achieved a reference sample-independent determination of the tip stray field distribution by scanning an MFM tip over a single NV center embedded within a diamond solid immersion lens used as a quantum magnetic field sensor with quasi-atomic resolution Sakar et al. (2021b). The MFM tip used in this investigation was a low-moment type (MFM_LM, TipsNano), coated with 20 nm CoCr and having a nominal tip radius of 30 nm. The distance between the tip apex and the NV center was 80 nm, while the measurement step size was set to 100 nm. The collected data points were here fitted using a pseudo-pole tip model as described in Ref. Häberle et al., 2012. Note, that in Ref. Sakar et al., 2021b, the fitting involved a Voigt function. The resulting NV measured $B_{z}^{\text{NV}}$ distribution is illustrated by the red curve in Fig. 2f.
@@
The same MFM tip was subsequently characterized by a quantitative magnetic force microscopy (qMFM) calibration procedure at the same lift height of 80 nm, employing a CoPt multilayer stack and the deconvolution process outlined in Eq. 2. This analysis yielded a reconstructed tip transfer function $B_{z}^{\text{qMFM}}$, represented by the black curve in Fig. 2f.
@@
The first notable observation in Fig. 2f is the significant discrepancy between $B_{z}^{\text{NV}}$ and $B_{z}^{\text{qMFM}}$ regarding both amplitude and width of the stray field peak. This difference translates into a distinct divergence in their respective Fourier spectra, as depicted in Fig. 3a. Specifically, the $B_{z}^{\text{qMFM}}$ distribution is devoid of the low-frequency components that are predominant in determining the peak width.
@@
From the tip stray field distribution, the instrument calibration function (ICF) can be derived as discussed in Eq. I. Once the ICF is determined, the effective magnetic surface charge density $\sigma_{\text{eff}}^{\text{sample}}$ of any SUT can be calculated from the MFM-measured phase shift data $\Delta\phi^{\text{SUT}}$ through deconvolution. This process again requires regularization, executed as outlined in Eq. 2, using an inverse Wiener filter with a regularization parameter $\alpha_{2}$:
@@
The $B_{z}^{\text{NV}}$ and $B_{z}^{\text{qMFM}}$ data from the specific MFM tip used in Ref. Sakar et al., 2021b were utilized to calculate ICFs and subsequently reconstruct the stray field distribution of a CoPt multilayer reference sample using the MFM phase shift data measured with this MFM tip. Despite the pronounced differences between $B_{z}^{\text{NV}}$ and $B_{z}^{\text{qMFM}}$, the reconstructed stray field distributions showed good agreement, as illustrated in Fig. 4 of Ref. Sakar et al., 2021b. However, the reasons for this good agreement and its limitations have not been discussed in detail, yet.
In the following, we will discuss the origins of the discrepancy between the real and reconstructed stray field distributions of the tip and its impact on qMFM measurements. We will generate and analyze simulated MFM images of samples with well-defined features in Section II. This analysis will allow us, in Section III, to explore how reference sample characteristics and measurement parameters influence the derived TTF. Additionally, we will discuss the applicability of calibration based on different types of reference samples for calibrated measurements of various SUTs in Section IV.
II Methods
@@
To investigate the discrepancies between qMFM-derived and NV-derived stray field distributions, we generate artificial MFM images using forward simulations. These simulations utilize generic reference structures with well-defined characteristic features. In all cases, the image sizes $\Delta_{x,y}$, pixel sizes $\delta_{x,y}$, and resolution resx,y are equal in x- and y-directions. For easier reading, an image size of 5.12 µm × 5.12 µm will be referred to as an image size of 5.12 µm. The same applies to the pixel size and resolution. We start from a typical measurement of a CoPt multilayer reference sample featuring a maze domain pattern. The pixel size was 10 nm, and the image size was 5.12 µm. Furthermore, we assume that the $B_{z}^{\text{NV}}$ data closely approximate the physical stray field distribution and thus can be considered as the “real” stray field distribution.
@@
Initially, we derive a binary magnetization pattern from the measured CoPt phase shift data (see Fig. 1). Using this pattern, we calculate the effective charge density using well-established magnetic parameters of the CoPt stack (Fig. 2a). To predict the expected MFM measurement response for the sample, we convolve the ICF, derived from $B_{z}^{\text{NV}}$ data (Fig. 2b), with the calculated effective charge density pattern. Additionally, a typical noise pattern is incorporated into the simulation (Fig. 2c). This noise pattern is acquired from an ambient condition MFM signal measured over a non-magnetic sample using a commercial MFM. An example of such a derived artificial MFM phase shift distribution is depicted in Fig. 2d, with a cross-section illustrated in Fig. 2e. A cross section of $B_{z}^{\text{NV}}$ (shown in red) is compared alongside the calculated $B_{z}^{\text{qMFM}}$ (shown in black) in Fig. 2f.
@@
Figure 2: Description of forward MFM simulation resulting in (d) by convolving the field of a simulated sample (a) with the tip’s field obtained from an NV measurement (b), and incorporating artificial noise (c). A line plot of the resulting simulated MFM signal is shown in (e). The comparison between $\mu_{0}H_{z}^{\text{tip}}$ measured by NV magnetometry and qMFM calibration is illustrated in (f).
@@
To assess the influence of different magnetic feature sizes in reference samples and varying measurement conditions, we also generate artificial reference samples. We begin with the original reference pattern image, which had 512 pixels, an image size of 5.12 µm and consequently a pixel size of 10 nm. We scale the original pattern to three larger sizes (10.24 µm, 15.36 µm and 20.48 µm), creating three additional reference patterns with domain sizes larger than the original by factors of x2, x3, and x4, respectively. These are marked green in Table 1. Due to this process, the pixel size increases to 20 nm (x2), 30 nm (x3) and 40 nm (x3), respectively. To emulate measurements with varying resolution, we reduce the pixel size of all three previously discussed patterns to 10 nm by a numerical interpolation. This process increases the number of pixels and thus changes the resolution (marked orange in Table 1). Finally, we cut out a 5.12 µm area from each of the latter high-resolution images, changing the image size while keeping domain size and pixel size (10 nm) constant (marked blue in Table 1). The latter images are visualized by the insets of Fig. 3b, c, and d. The inset in Fig. 3a shows the original reference sample. We thus generated a set of artificial reference sample patterns with different domain sizes, resolution, and image sizes which will be used to analyze the impact of these parameters on the calibration process. The Fourier spectra of these reference patterns are shown in the left panels of Fig. 3, together with the spectrum of the original reference pattern in light red. Additionally, the Fourier spectrum of the tip’s stray field distribution is plotted in grey, derived from the NV-measured stray field distribution of the tip.
@@
Using these datasets, we calculate emulated MFM data as described in the previous section following the approach depicted in Fig. 2a-d. Subsequently, the TTF is derived from each of the simulated noisy MFM images using the standardized approach. This involves discrimination, calculation of the effective charge pattern, and regularized deconvolution employing a Wiener filter (see Ref. IEC, 2021). In all cases, the regularization parameter for the Wiener filter is determined by applying the L-curve criterion to derive the optimal regularization and avoid ambiguities (for detailed methodology, see Ref. IEC, 2021).
These simulated results are discussed in the following Section III.
III Influence of the Reference Sample on the Derived qMFM Tip Transfer Function

### v2

@@ Abstract
— Copyright 2025 Author(s). This article is distributed under a Creative Commons Attribution-NonCommercial- NoDerivs 4.0 International (CC BY-NC-ND) License.
— This is the author’s peer reviewed, accepted manuscript. However, the online version of record will be different from this version once it has been copyedited and typeset. PLEASE CITE THIS ARTICLE AS DOI: 10.1063/5.0288740
— This article may be downloaded for personal use only. Any other use requires prior permission of the author and AIP Publishing. This article appeared in Review of Scientific Instruments 96, 123708 (2025) and may be found at https://doi.org/10.1063/5.0288740.
@@ I INTRODUCTION
In common two-pass mode MFM, a magnetically coated tip mounted on an oscillating cantilever is scanned over the sample at a measurement height $z$. The phase shift $\Delta\phi$ of the oscillating cantilever is monitored to measure the interaction strength between the tip and the sample. For many applications, it is essential to obtain not only qualitative data but also quantitative data in terms of the sample’s stray field distribution using either the magnetic field $H$ (in units of A/m) or the magnetic flux density $B=\mu_{0}H$ (in units of Tesla). Such quantitative MFM (qMFM) requires calibrating the MFM system, that means finding the functional relationship that connects the magnetic field $B^{\text{sample}}(x,y)$ of a thin film sample (which can be modeled as magnetic surface charges projected onto the upper surface of the sample) to the measured MFM signal, thereby providing spatially resolved phase shift data $\Delta\phi(x,y)$.
@@
Initially, calibrations relied on interpreting the magnetically coated etched silicon tips as point-like dipoles or monopoles interacting with the local stray field, fitting this model to reference measurements. While this approach is justified for nanotube-shaped tips, when working with coated tips, these models exhibit significant limitations due to their strong dependence on the feature sizes of the sample Vock2010; McVitie2001; Schendel2000. These limitations can be overcome by determining the spatial frequency-dependent transfer function of the MFM system, known as the instrument calibration function (ICF), as introduced by Hug et al. Hug1998 and Schendel et al. Schendel2000.
To obtain the ICF, it is necessary to determine the magnetic stray field of the tip. In this paper, we demonstrate that while a high-quality reconstruction of the tip’s stray field in real space appears desirable, it is more critical to ensure a strong overlap of frequency components between the reference sample and the SUT than to achieve an accurate real-space reconstruction of the stray field distribution.
The paper is structured as follows: In Section II, we introduce the basic theory of MFM calibration and compare a reference sample-based tip stray field calibration with an NV-based tip stray field calibration. Following this, we discuss the origins of the discrepancies between the real and reconstructed stray field distributions of the tip and their impact on qMFM measurements in Section III. To achieve this, we generate and analyze simulated MFM images of samples with well-defined features. This analysis will allow us, in Section IV, to explore how the characteristics of the reference sample and measurement parameters influence the derived TTF. Additionally, we discuss the applicability of calibration based on different types of reference samples for calibrated measurements of various SUTs in Section V, showcasing the impact of (lacking) k-space overlap.
II MFM Tip Calibration
Instrument Calibration Function and Tip Transfer Function
The ICF is a function of the cantilever’s mechanical properties and of the gradient of the force exerted on the tip. This force arises from the non-local interaction between the vertical component of the spatially extended MFM tip stray field distribution $B_{z}^{\text{tip}}$ and the magnetization of the sample under test (SUT), as illustrated in Fig. 1a. The spatially extended tip stray field distribution at the sample surface thus governs the point-spread function character of the ICF.
Figure 1: (a) MFM measurements represented as the convolution of the stray field from the tip with the sample magnetization. (b) MFM phase shift data for a 5.12 µm x 5.12 µm scan of a CoPt multilayer stack, displaying a maze domain pattern. A cross-section along white dashed is shown in black color. The corresponding wavelength density spectrum in Fourier space is shown in red color.
@@ II MFM Tip Calibration
Mathematically, the measured cantilever phase shift can be conveniently described in partial Fourier space $(x,y,z)\to(k_{x},k_{y},z)=(\mathbf{k},z)$, where $k=\sqrt{k_{x}^{2}+k_{y}^{2}}$. In this space, the interaction between the tip’s stray field and the sample’s stray field can be represented by the convolution of the tip’s stray field at the sample surface with the effective magnetic surface charge density Corte‐Leon2020 of the sample, $\sigma_{\text{eff}}^{\text{sample}}$, which translates to a multiplication:
$\Delta\phi(\mathbf{k},z)=$ $\text{ICF}(\mathbf{k},z)\cdot\sigma_{\text{eff}}^{\text{sample}}$
@@
Here, $Q$ and $C$ represent the quality factor and the stiffness of the oscillating cantilever, respectively. The lever correction function (LCF) accounts for the finite oscillation amplitude $A$ and the canting angle $\Theta$ of the cantilever. Consequently, for a quantitative analysis of a SUT, it is essential to know both the shape and amplitude of the tip stray field distribution, as well as the scalar geometrical and mechanical parameters of the cantilever. This information allows for the calculation of the effective magnetic charge density distribution of any sample - and thus its stray field - by deconvolving the measured phase shift signal. Therefore, calibrating MFM requires determining $B_{z}^{\text{tip}}(k)$. By exploiting Eq. II, this can be achieved by measuring the phase shift distribution $\Delta\phi^{\text{ref}}$ of a reference sample with a well-known effective magnetic surface charge density $\sigma_{\text{eff}}^{\text{ref}}$.
Ideally, deconvolution to obtain $B_{z}^{\text{tip}}$ from Eq. II can be described as a division in Fourier space. However, due to measurement noise, this deconvolution is ill-posed and necessitates regularization, a method used to impose additional constraints to stabilize the inversion process and reduce noise in the reconstructed image. This is conveniently performed using an inverse Wiener filter (see Ref. Wiener1949) with a regularization parameter $\alpha_{1}$ (Eq. 2) as discussed in the relevant IEC standard on quantitative MFM IEC2021. Typically, $\alpha_{1}$ is determined by an L-curve criterion Hu2020. Due to regularization, the reconstructed tip stray field distribution becomes frequency-filtered. To distinguish the reconstructed $B_{z}^{\text{tip}}$ from the actual $B_{z}^{\text{tip}}$, it is often referred to as the tip transfer function (TTF).
@@
Another factor influencing the TTF is the choice of the reference sample. To date, only stripe domain patterns have been tested and validated as suitable reference samples Hug1998; Hu2020; Sakar2021. These samples are characterized by distinct domain patterns with specific domain sizes and narrow transitions between domains.
It is crucial to note that the calibration principle described here is based on the premise that neither the tip nor the reference sample is altered during measurements. Notably, the calibration sample must be measured in advance and subsequently re-measured after a SUT measurement to rule out any potential tip damage during the SUT measurement IEC2021. A calibration is only applicable if these reference sample measurements coincide.
Tip Calibration using a Reference Sample
An MFM image of a typical stripe domain sample (here a maze domain pattern that is composed of complex, interconnected pathways resembling a maze), specifically a CoPt multilayer stack as referenced in Ref. Hu2020, is illustrated in Fig. 1b alongside a cross-section plot and its circularly averaged Fourier spectrum.
Determining the underlying magnetization pattern, which is necessary for calculating the effective surface charge density, involves discriminating the MFM image into up- and down-magnetized domains followed by applying a domain wall operator. The thereof calculated effective charge density pattern provides a good approximation of the real density pattern (see Fig. 1) and is subsequently used in the deconvolution process.
@@
Although the fundamental validity of transfer function-based calibration has been demonstrated for these specific samples, its limitations - stemming from dependencies on the reference sample characteristics and measurement parameters such as image size and pixel resolution - have not yet been thoroughly investigated.
@@
Additionally, understanding the relationship between the TTF and the actual tip stray field distribution is crucial but challenging due to difficulties in directly assessing the tip stray field distribution with sufficient resolution and sensitivity. Over the past three decades, only a few attempts have been published that characterize the magnetic stray-field distribution of MFM tips using techniques such as the Hall effect Abelmann1998; Vock2014, Lorentz tomography or holography Dai2018; Neu2018; Gao2004. Recent advancements in NV microscopy (see Ref. Casola2018) have introduced new methods for highly spatially resolved characterization of magnetic stray fields.
NV-measured Tip Stray Field Distribution
@@
Exploiting these techniques, a recent study achieved a reference sample-independent determination of the tip stray field distribution by scanning an MFM tip over a single NV center embedded within a diamond solid immersion lens used as a quantum magnetic field sensor with quasi-atomic resolution Sakar2021a. The MFM tip used in this investigation was a low-moment type (MFM_LM, TipsNano), coated with 20 nm CoCr and having a nominal tip radius of 30 nm. The distance between the tip apex and the NV center was 80 nm, while the measurement step size was set to 100 nm. The collected data points were here fitted using a pseudo-pole tip model as described in Ref. Haeberle2012. Note, that in Ref. Sakar2021a, the fitting involved a Voigt function. The resulting NV measured $B_{z}^{\text{NV}}$ distribution is illustrated by the red curve in Fig. 2.
@@
The same MFM tip was subsequently characterized by a quantitative magnetic force microscopy (qMFM) calibration procedure at the same lift height of 80 nm, employing a CoPt multilayer stack and the deconvolution process outlined in Eq. 2. This analysis yielded a reconstructed tip transfer function $B_{z}^{\text{qMFM}}$, represented by the black curve in Fig. 2.
As introduced in Section II, it is essential to verify that no tip-sample interaction has altered the qMFM calibration. This verification was achieved by conducting repeated measurements with different tips in a NV $\rightarrow$ MFM $\rightarrow$ NV $\rightarrow$ MFM sequence, which yielded consistent results, indicating that no tip damage occurred. More information can be found in the supplementary materials of [Sakar2021a]. Fig. 2 presents one measurement from the verified series.
Figure 2: Comparison between $\mu_{0}H_{z}^{\text{tip}}$ measured by NV magnetometry (red) and qMFM calibration (black). The line profiles display a cross-section through $\mu_{0}H_{z}^{\text{tip}}$ as indicated by the dashed line in the insets.
@@
The first notable observation in Fig. 2 is the significant discrepancy between $B_{z}^{\text{NV}}$ and $B_{z}^{\text{qMFM}}$ regarding both amplitude and width of the stray field peak. This difference translates into a distinct divergence in their respective Fourier spectra. Specifically, the $B_{z}^{\text{qMFM}}$ distribution is devoid of the low-frequency components that are predominant in determining the peak width.
@@
From the tip stray field distribution, the instrument calibration function (ICF) can be derived as discussed in Eq. II. Once the ICF is determined, the effective magnetic surface charge density $\sigma_{\text{eff}}^{\text{sample}}$ of any SUT can be calculated from the MFM-measured phase shift data $\Delta\phi^{\text{SUT}}$ through deconvolution. This process again requires regularization, executed as outlined in Eq. 2, using an inverse Wiener filter with a regularization parameter $\alpha_{2}$:
@@
The $B_{z}^{\text{NV}}$ and $B_{z}^{\text{qMFM}}$ data from the specific MFM tip used in Ref. Sakar2021a were utilized to calculate ICFs and subsequently reconstruct the stray field distribution of a CoPt multilayer reference sample using the MFM phase shift data measured with this MFM tip. Despite the pronounced differences between $B_{z}^{\text{NV}}$ and $B_{z}^{\text{qMFM}}$, the reconstructed stray field distributions showed good agreement, as illustrated in Fig. 4 of Ref. Sakar2021a. However, the reasons for this good agreement and its limitations have not been discussed in detail, yet.
III Methods
@@
To investigate the discrepancies between qMFM-derived and NV-derived stray field distributions, we generate artificial MFM images using forward simulations. These simulations utilize generic reference structures with well-defined characteristic features. In all cases, the image sizes $\Delta_{x,y}$, pixel sizes $\delta_{x,y}$, and resolution resx,y are equal in x- and y-directions. For easier reading, an image size of 5.12 µm × 5.12 µm will be referred to as an image size of 5.12 µm. The same applies to the pixel size and resolution. We start from a typical measurement of a CoPt multilayer reference sample featuring a maze domain pattern. The pixel size was 10 nm, and the image size was 5.12 µm. Furthermore, we assume that the $B_{z}^{\text{NV}}$ data closely approximate the physical stray field distribution and thus can be considered as the "real" stray field distribution.
@@
Initially, we derive a binary magnetization pattern from the measured CoPt phase shift data (see Fig. 1). Using this pattern, we calculate the effective charge density using well-established magnetic parameters of the CoPt stack. The layer architecture of the CoPt stack is as follows: Pt(2 nm)/[(Co(0.4 nm)/Pt(0.9 nm)]100/Pt(5 nm)/Ta (5 nm)/SiOx/Si(100), exhibiting a saturation magnetization of $M_{s}=(500\pm 30)$ kA/m, a Bloch-type domain transition with a domain wall width of approximately $\delta_{DW}=16$ nm, and an average domain width of about 170 nm Hu2020; Vock2014. Using these values, Fig. 3a can be derived. To predict the expected MFM measurement response for the sample, we convolve the ICF, derived from $B_{z}^{\text{NV}}$ data (Fig. 3b), with the calculated effective charge density pattern. Additionally, a typical noise pattern is incorporated into the simulation (Fig. 3c). This noise pattern is acquired from an ambient condition MFM signal measured over a non-magnetic sample using a commercial MFM. An example of such a derived artificial MFM phase shift distribution is depicted in Fig. 3d, with a cross-section illustrated in Fig. 3e.
@@
Figure 3: Description of forward MFM simulation resulting in (d) by convolving the field of a simulated sample (a) with the tip’s field obtained from an NV measurement (b), and incorporating artificial noise (c). A line plot (along the dashed line) of the resulting simulated MFM signal is shown in (e).
@@
To assess the influence of different magnetic feature sizes in reference samples and varying measurement conditions, we generate artificial reference samples, beginning with the original reference pattern image, which had 512 pixels, an image size of 5.12 µm and consequently a pixel size of 10 nm. We scale the original pattern to three larger sizes (10.24 µm, 15.36 µm and 20.48 µm), creating three additional reference patterns with domain sizes larger than the original by factors of x2, x3, and x4, respectively. These are marked green in Table 1. Due to this process, the pixel size increases to 20 nm (x2), 30 nm (x3) and 40 nm (x4), respectively. To emulate measurements with varying resolution, we reduce the pixel size of all three previously discussed patterns to 10 nm by a numerical interpolation. This process increases the number of pixels and thus changes the resolution (marked orange in Table 1). Finally, we cut out a 5.12 µm area from each of the latter high-resolution images, changing the image size while keeping domain size and pixel size (10 nm) constant (marked blue in Table 1).
We thus generated a set of artificial reference sample patterns with different domain sizes, resolution, and image sizes which will be used to analyze the impact of these parameters on the calibration process. Using these datasets, we calculate emulated MFM data as described in the previous section following the approach depicted in Fig. 3a-d. Subsequently, the TTF is derived from each of the simulated noisy MFM images using the standardized approach. This involves discrimination, calculation of the effective charge pattern, and regularized deconvolution employing a Wiener filter (see Ref. IEC2021). In all cases, the regularization parameter for the Wiener filter is determined by applying the L-curve criterion to derive the optimal regularization and avoid ambiguities (for detailed methodology, see Ref. IEC2021). These simulated results are discussed in the following Section IV.
@@
IV Influence of the Reference Sample on the Derived qMFM Tip Transfer Function